OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!!!!!!! A rectangular fountain has an area of (16x^2+8x+1)ft squared. The dimensions of the rectangle have the form ax+b, where a and b are whole numbers. Write an expression for the perimeter of the fountain. Then find the perimeter when x=2 feet.

OpenStudy (anonymous):

I KNOW YOU ARE SUPPOSED TO FACTOR BUT DON'T KNOW HOW???

OpenStudy (anonymous):

P=4(4x+1) i think so then 36

OpenStudy (anonymous):

could you please show steps

OpenStudy (anonymous):

16x^2 + 4x +4x +1 4x(4x +1)+1(4x+1) (4x +1)^2 = area 4(4x=1) =P

OpenStudy (anonymous):

thanks could you help with more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1 min please

OpenStudy (anonymous):

A square tabletop has an area of (9x^2-90x+225) cm^2. The dimensions of the tabletop have the form cx-d, where c and d are whole numbers. Write an expression for the perimeter when x=25 cm.

OpenStudy (anonymous):

one sec, i'm working it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u get how to factor it? by browser keeps stopping so i keep losing it

OpenStudy (anonymous):

kind of.....

OpenStudy (anonymous):

P = 4(9x - 45) 36x -180 36(25)-180 720 cm

OpenStudy (anonymous):

how did you get 4(9x-45) please show steps

OpenStudy (anonymous):

9x^2 -45x -45x + 225 9x(x-5)-45(x-5) (9x-45)(x-5) 9(x-5)(x-5) 9(x-5)^2 = area, so one side=9(x-5)